Elastic Electroproduction of ρ Mesons at HERA

H1 Collaboration

DESY, Hamburg, Germany

The elastic electroproduction of ρ mesons is studied at HERA with the H1 detector for a photon virtuality in the range 1 < Q^{2} < 60 GeV^{2} and for a hadronic centre of mass energy in the range 30 < W < 140 GeV. The shape of the (ππ) mass distribution in the ρ resonance region is measured as a function of Q^{2}. The full set of ρ spin density matrix elements is determined, and evidence is found for a helicity flip amplitude at the level of 8 ± 3% of the non-flip amplitudes. Measurements are presented of the dependence of the cross section on Q^{2}, W and t (the four-momentum transfer squared to the proton). They suggest that, especially at large Q^{2}, the γ^{∗}p cross section develops a stronger W dependence than that expected from the behaviour of elastic and total hadron−hadron cross sections.

References

1. H1 Collaboration, C. Adloff et al, DESY-99-010 / submitted to Eur.Phys.J. C